OpenStudy (shelby1290):

Find the equation of the line. parallel to the x-axis, passing through (3,-6)

OpenStudy (anonymous):

it is parallel to the x-axis, so it will look like this:  the \(line\ will\ always\ pass\ through\ \large{y=-6}\), since it is parallel to the x-axis. therefore the equation will be..?

OpenStudy (shelby1290):

wait

OpenStudy (shelby1290):

it also needs to include the formula: y=m(x-p)+q

OpenStudy (shelby1290):

sorry i forgot to mention that...

OpenStudy (shelby1290):

and by "it" i mean like the whole question

OpenStudy (anonymous):

parrallel to x axis means line has slope =0

OpenStudy (anonymous):

since it is \(parallel\ to\ the\ x-axis\), the slope(m) will be 0 \(y= 0(x-p)+q\\y=0(x-3)-6\\ y= ? \)

OpenStudy (anonymous):

so standard equation y=mx+c with slope =0 y=c

OpenStudy (anonymous):

so put the given point in equation as no x term in equation ..so just plugged in the y value and calculate value of C

OpenStudy (anonymous):

-6=c so c=-6 as y=c=-6 so y=-6

OpenStudy (shelby1290):

ok what i've picked up is.. m=0 (p,q)=(3,-6) y=0(x-3)-6

OpenStudy (shelby1290):

whats next..?

OpenStudy (shelby1290):

0x-3-6 right?

OpenStudy (anonymous):

loll 0*x-0*3-6

OpenStudy (anonymous):

y=0-0-6

OpenStudy (anonymous):